# LLMs Do Not Have Human-Like Working Memory

### v1

@@ Abstract
Human working memory is an active cognitive system that enables not only the temporary storage of information but also its processing and utilization. Without working memory, individuals may produce unreal conversations, exhibit self-contradiction, and struggle with tasks requiring mental reasoning. In this paper, we demonstrate that Large Language Models (LLMs) lack this human-like cognitive ability, posing a significant challenge to achieving artificial general intelligence. We validate this claim through three experiments: (1) Number Guessing Game, (2) Yes or No Game, and (3) Math Magic. Experimental results on several model families indicate that current LLMs fail to exhibit human-like cognitive behaviors in these scenarios. By highlighting this limitation, we aim to encourage further research in developing LLMs with improved working memory capabilities.
Large Language Model,Working Memory
@@ 1 Introduction
Imagine the following scenario: You select a number between one and ten. When ready, you are asked, “Is the number greater than five?” Upon answering, others can infer that the number has likely [^1] A possibility remains that your response was made by chance. entered your conscious awareness, as its clear perception is necessary for making the comparison and providing a response.
Figure 1: When ChatGPT says it already has a number in mind, and it is not 4, how can we know whether ChatGPT is lying?
Such conscious awareness is commonly referred to as Working Memory (Atkinson & Shiffrin, 1968). In contrast to long-term memory, working memory is the system required to maintain and manipulate information during complex tasks such as reasoning, comprehension, and learning (Baddeley, 2010). Deficits in working memory can impair information processing and hinder effective communication (Gruszka & Nkecka, 2017; Cowan, 2014).
In this paper, we demonstrate that, despite exhibiting increasingly human-like abilities (Huang et al., 2024b, a), Large Language Models (LLMs) lack a fundamental aspect of human cognition: working memory. As a result, LLMs generate unrealistic responses, display self-contradictions, and fail in tasks requiring mental manipulation.
Previous studies have used the N-back task (Kirchner, 1958) to evaluate the working memory of LLMs (Gong et al., 2024; Zhang et al., 2024). However, a fundamental limitation of these tests is that the critical information needed to answer correctly is in the input context accessible to LLMs. This differs markedly from human testing, where participants cannot view prior steps. LLMs, in contrast, can simply attend to earlier input tokens retained in their context window. To more accurately assess working memory, it is essential to design experiments in which the key information is not explicitly present in the context.
@@
The core challenge is to prove that there is actually nothing in LLMs’ mind without knowing what exactly is in their mind. To investigate this limitation, we design three experiments—(1) Number Guessing Game, (2) Yes-No Game, and (3) Math Magic—that test whether LLMs can maintain information internally without explicitly externalizing it in the context. Experimental results reveal that LLMs consistently fail in this capacity, regardless of model family (GPT (Hurst et al., 2024), Qwen (Yang et al., 2024), DeepSeek (Liu et al., 2024; Guo et al., 2025), LLaMA (Grattafiori et al., 2024)) or reasoning approach (Chain-of-Thought (CoT) (Wei et al., 2022) or o1-like reasoning (Jaech et al., 2024)).
2 Number Guessing Game
@@
Findings.
Fig. 2 presents the probabilities of “Yes” responses in each model for numbers from one to ten. Two key observations emerge: (1) Most LLMs never produce a “Yes” response; “No” dominates across models. This suggests that LLMs generate unrealistic distributions, implying an inability to produce realistic responses. Given that LLMs generally follow instructions and do not deliberately deceive, we attribute this behavior to their failure to internally “think of” a number. (2) When LLMs do respond affirmatively, they exhibit a marked preference for the number seven. This tendency mirrors patterns observed in humans (Miller, 1956; Kubovy & Psotka, 1976).
@@
Table 2: The sum of probabilities of GPT-4o-2024-08-06 responding “Yes” for all numbers in different ranges. Color intensity reflects proximity to one: red indicates values closer to zero, while blue signifies values greater than one.
@@ 3 Yes-No Game
Table 4: Count of failures of GPT-4o-Mini-2024-07-18 and GPT-4o-2024-08-06 on different properties.
@@ 4 Math Magic
User Take the first {random_number_a} numbers and move them to the end.
@@
Assistant Okay! Now I have {8 - random_number_c} numbers. What’s next?
User Take the first number and move it to the end. Repeat this seven times.
@@
User Take the first number and move it to the end, then remove the second number. Repeat this {7 - random_number_c} times.
@@
Since the 1089 game is well known and likely included in the training data of LLMs, our preliminary experiments show that LLMs can recognize the game and thus invalidate the evaluation. Therefore, to more effectively assess LLMs’ multi-step mental manipulation, we select a more complicated math magicbased on the Josephus Problem (Schumer, 2002). In this task, participants are asked to imagine four numbers and perform a sequence of operations, including duplication, rotation, and removal. The procedure is shown in Table 5. Ultimately, only two numbers remain, and mathematical constraints guarantee they are identical. In our experiment, LLMs are prompted to secretly think of four numbers and mentally execute the sequence of randomized operations. We measure the proportion of trials (out of 150) in which the LLM completes the task correctly.
Table 6: Model performance on the Math Magic. GPT-4.1-2025-04-14 fails to complete the task, incorrectly assuming that the necessary numerical inputs are missing.
@@
LLaMA-3.3-70B-Instruct-Turbo 21/150 14.0
LLaMA-3.1-8B-Instruct-Turbo 33/150 22.0
LLaMA-3.1-70B-Instruct-Turbo 12/150 8.0
LLaMA-3.1-405B-Instruct-Turbo 28/150 18.7
Qwen2.5-7B-Instruct-Turbo 0/150 0.0
Qwen2.5-72B-Instruct-Turbo 8/150 5.3
DeepSeek-V3 0/150 0.0
Table 7: Performance of LLMs with CoT or LRMs on the Math Magic. Similar to GPT-4.1-2025-04-14, o4-Mini-2025-04-16 also fails to complete the task.
@@
GPT-4o-Mini-2024-07-18 3/150 2.0
GPT-4o-2024-08-06 32/150 21.3
LLaMA-3.3-70B-Instruct-Turbo 33/150 22.0
Qwen2.5-72B-Instruct-Turbo 23/150 15.3
DeepSeek-V3 18/150 12.0
o1-Mini-2024-09-12 25/150 16.7
o3-Mini-2025-01-31 3/150 2.0
o4-Mini-2025-04-16 - -
QwQ-32B 45/150 30.0
DeepSeek-R1 59/150 39.3
@@
Table 6 reports the accuracy of prompting models to output the two numbers directly. Most LLMs perform poorly on this task, with notably higher accuracy observed in the LLaMA model family. This finding aligns with results from the number guessing game shown in Table 1, where LLaMA models generate more realistic distributions than other models. Together, these consistent results suggest that current LLMs generally lack the working memory required for such tasks.
We also investigate whether CoT prompting can enhance performance on this task. Table 7 presents the results of prompting models to reason step by step, as well as the performance of o1-like long reasoning models (LRMs). The best-performing model is DeepSeek-R1, achieving an accuracy of 39.3%. Most models attain only 10–20% accuracy—marginally above the random guessing baseline of 10%—highlighting substantial room for improvement. Notably, as observed in the number guessing experiment, models exhibit a preference for the number seven, which appears to influence performance here. Of the 59 correct predictions by DeepSeek-R1, 48 (81.4%) involve selecting the number 7; for o1-Mini, the number is 22 out of 25 (88%). These results suggest that models may not fully comprehend the task and instead rely on biased number preferences.
In conclusion, current LLMs struggle with a class of tasks that remain challenging even with CoT prompting. These tasks demand internal object imagination and mental manipulation, relying on an effective working memory.
5 Related Work
Working memory has been widely discussed in the context of LLMs. However, most studies use the term to describe LLMs’ capacity to process information within a fixed context window (Li et al., 2023; Guo et al., 2023). Gong et al. (2024); Zhang et al. (2024) evaluate LLM working memory using N-back tasks. As noted in §1, however, such human-designed tests may not be valid for LLMs.
A separate line of work defines LLM working memory as an explicitly designed external module for LLM agents (Hu et al., 2024; Zeng et al., 2024). Wang et al. (2024) incorporate working memory to enhance reasoning, while Kang et al. (2024) use it to improve training efficiency. These approaches, however, do not explore whether LLMs possess a form of human-like working memory—such as the ability to mentally simulate and manipulate objects.

### v2

# Language Models Do Not Have Human-Like Working Memory

@@ Abstract
While Large Language Models (LLMs) exhibit remarkable reasoning abilities, we demonstrate that they fundamentally lack a core aspect of human cognition: working memory. Human working memory is an active cognitive system that enables not only the temporary storage of information but also its processing and utilization. Without working memory, individuals may produce unreal conversations, exhibit self-contradiction, and struggle with tasks requiring mental reasoning. Existing evaluations using N-back or context-dependent tasks fail as they allow LLMs to exploit accessible context rather than retain latent information. We introduce three novel tasks—(1) Number Guessing, (2) Yes-No Deduction, and Math Magic—that isolate internal representation from external context. Across seventeen frontier models spanning four major model families, we consistently observe irrational or contradictory behaviors, highlighting LLMs’ inability to retain and manipulate latent information. Our work establishes a new benchmark for evaluating working memory in LLMs and identifies this deficit as a critical obstacle to artificial general intelligence. Code and prompts for the experiments are available at https://github.com/penguinnnnn/LLM-Working-Memory.
Figure 1: When LLMs say they already have a number in mind, and it is not 4, how can we know whether LLMs are lying, or even thinking of nothing?
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated strong performance across diverse domains, including legal advice Guha et al. (2023), education Wen et al. (2024), translation Jiao et al. (2023), and healthcare Yang et al. (2024b). Beyond task performance, recent studies have examined their intrinsic behaviors, often drawing parallels with human cognition, such as personality Huang et al. (2024b), emotion Huang et al. (2024a), empathy Sorin et al. (2024), theory-of-mind Liu et al. (2024b), and values Wang et al. (2024b).
One such behavior is memory, which has attracted increasing attention from both industry and research communities. OpenAI was the first to introduce a memory module in ChatGPT (February 2024). [^1] https://openai.com/index/memory-and-new-controls-for-chatgpt/ By mid-2025, xAI, Anthropic, and Google integrated memory into Grok, [^2] https://x.com/grok/status/1912670182012801156 Claude, [^3] https://www.anthropic.com/news/claude-4 and Gemini, [^4] https://blog.google/products/gemini/temporary-chats-privacy-controls/ respectively. Sam Altman even emphasized that “people want memory,” identifying it as a central feature of the forthcoming GPT-6.555https://futurism.com/disastrous-gpt-5-sam-altman-hyping-up-gpt-6 On the research side, memory modules are also actively studied Zhang et al. (2025); Hatalis et al. (2023); Xu et al. (2025).
What is memory? Atkinson & Shiffrin (1968) categorized memory by retention timescale: sensory (1ms–2s), short-term (seconds), and long-term (hours to lifetime). Long-term memory enables information storage over extended periods, whereas short-term, or Working Memory, maintains and manipulates information during complex tasks such as reasoning, comprehension, and learning Baddeley & Hitch (1974). These distinctions have also been adopted in machine learning, corresponding to representational learning for raw inputs (sensory memory), in-context computation at test time (working memory), and access to external databases (long-term memory) Weng (2023). As a central aspect of human cognition, memory has been widely benchmarked Du et al. (2025). However, working memory remains relatively underexplored.
Existing studies often adopt the N-back task Kirchner (1958) to assess LLMs’ working memory Gong et al. (2024); Zhang et al. (2024). Yet, a fundamental limitation arises: the critical information for correct responses remains accessible in the model’s input context. Unlike humans, who cannot revisit prior steps, LLMs can simply attend to earlier tokens within their context window. Fig. 2 illustrates this discrepancy. To more faithfully evaluate working memory, it is necessary to design experiments where the key information is not explicitly present in the context. Imagine the following scenario: You select a number between one and ten. When ready, you are asked, “Is the number greater than five?” If you answer, observers can reasonably infer that the number has entered your conscious awareness (i.e., your working memory), since clear perception is necessary to perform the comparison and provide a response. [^6] A possibility remains that the response was given by chance.
We ask: Do LLMs possess human-like working memory beyond evaluations that merely test their ability to attend to visible information? AI models are often viewed as reasoning in two modes: (1) the token space over sequences Wei et al. (2022); Yao et al. (2023), and (2) the latent space over activations Hao et al. (2024); Geiping et al. (2025). Our focus is on the second mode, where evaluating working memory provides insight into whether models can hold and manipulate latent concepts without explicit externalization. Success in this capacity could enhance reasoning without reliance on Chain-of-Thought (CoT), as it directly tests the model’s ability to maintain objects and concepts internally. Conversely, deficits in working memory impair information processing in humans Gruszka & Nkecka (2017); Cowan (2014), and in LLMs manifest as unrealistic outputs, self-contradictions, and failures on tasks requiring mental manipulation.
The central challenge in designing such evaluations is epistemic: how can we demonstrate the absence of internal memory when we cannot directly observe a model’s mind? To address this limitation, we design three experiments—(1) Number Guessing, (2) Yes–No Deduction, and (3) Math Magic—that test whether LLMs can internally maintain information that is not explicitly present in the input context. Our experiments span 17 state-of-the-art LLMs, both proprietary and open-source, and reveal a consistent inability to perform such tasks. The failure is uniform across model families (GPT Hurst et al. (2024), Qwen Yang et al. (2024a), DeepSeek Liu et al. (2024a); Guo et al. (2025), LLaMA Grattafiori et al. (2024)) and reasoning approaches (CoT Wei et al. (2022) or o1-like reasoning Jaech et al. (2024)).
2 Background
Figure 2: a. An illustration of how “N-Back” tasks are performed. b. Humans see the stimuli one after one, forcing them to put the information in working memory. c. Researchers put all stimuli into context, enabling LLMs to easily find the answers.
2.1 How Working Memory is Evaluated
Human working memory is typically assessed using behavioral paradigms that require individuals to maintain and update information over short intervals. Common examples include the digit span task Miller (1956), where participants recall sequences of numbers of increasing length, and the N-back task Kirchner (1958), which requires identifying whether the current stimulus matches one presented $N$ steps earlier. These tasks are widely used because they probe the ability to maintain and manipulate information that is no longer externally visible, thereby capturing the essence of working memory function.
In the context of LLMs, working memory has been widely discussed. However, most studies use the term to describe an LLM’s capacity to process information within a fixed context window Li et al. (2023); Guo et al. (2023). Gong et al. (2024); Zhang et al. (2024) evaluate LLMs using N-back tasks. As noted in §1, such human-designed tests are not directly valid for LLMs, since models can simply attend to retained tokens in their context window rather than actively maintaining information.
2.2 Existing Engineering Solutions
A separate line of work introduces explicit engineering solutions by equipping LLMs with external memory modules Hu et al. (2024); Zeng et al. (2024). For example, Wang et al. (2024a) incorporate symbolic working memory to enhance reasoning, while Kang et al. (2024) use it to improve training efficiency. Other approaches implement scratch spaces Lanchantin et al. (2023) or internal CoT mechanisms Jaech et al. (2024), which may partially mitigate working memory limitations identified in this paper. However, these methods do not address whether LLMs possess an intrinsic working memory capacity, analogous to humans’ ability to mentally simulate and manipulate objects. The distinction can be illustrated with an analogy: modern calculators make arithmetic trivial, yet schools continue to assess addition, subtraction, multiplication, and division. The purpose is not solely to solve problems efficiently, but to reveal individuals’ underlying cognitive abilities. Similarly, while engineering techniques can extend an LLM’s effective memory, we are ultimately interested in whether the model itself has the basic cognitive capability of working memory. Without it, a model may function adequately through external tools, but its intrinsic reasoning ability remains fundamentally limited.
3 Number Guessing
@@
Findings.
Fig. 3 presents the probabilities of “Yes” responses in each model for numbers from one to ten. Two key observations emerge: (1) Most LLMs never produce a “Yes” response; “No” dominates across models. This suggests that LLMs generate unrealistic distributions, implying an inability to produce realistic responses. Given that LLMs generally follow instructions and do not deliberately deceive, we attribute this behavior to their failure to internally “think of” a number. (2) When LLMs do respond affirmatively, they exhibit a marked preference for the number seven. This tendency mirrors patterns observed in humans Miller (1956); Kubovy & Psotka (1976).
Table 2: The sum of probabilities of GPT-4o-2024-08-06 responding “Yes” for all numbers in different ranges.
@@ 4 Yes-No Deduction
Setup.
Table 4: Count of failures of GPT-4o-Mini-2024-07-18 and GPT-4o-2024-08-06.
@@
Table 5: Ablation studies using GPT-4o-2024-08-06.
Model Failure V W L D H
Hint 194 37 39 60 37 21
All 158 18 29 21 55 35
Hint + All 145 13 32 34 46 20
Ablation.
To ensure that the observed failures are not simply due to LLMs’ inability to rank objects by the five properties, we conduct the following ablation studies: (1) Hint: At the beginning of the prompt, we provide GPT-4o with the object rankings defined in Table 3. (2) All: For each question, we specify the target object $\mathcal{O}$ by stating explicitly that “the object you are considering is $\mathcal{O}$.” (3) Hint + All: We combine the above two settings. Results are shown in Table 5. Two key findings emerge: (1) Even when given hints that guarantee knowledge of the correct ranking, GPT-4o still frequently fails, suggesting that the task critically depends on long-context reasoning. (2) When the object identity is provided explicitly, error rates decrease, likely because the long-context reasoning task is reduced to a short-context one.
@@ 5 Math Magic
User Move the first {random_number_a} numbers to the end.
@@
Assistant Okay! Now I have {7 - random_number_c} numbers. What’s next?
User Move the first number to the end. Repeat this seven times.
@@
User Remove the second number, and then move the first number to the end. Repeat this {6 - random_number_c} times.
@@
Since the 1089 game is well known and likely included in the training data of LLMs, our preliminary experiments show that LLMs can recognize the game and thus invalidate the evaluation. Therefore, to more effectively assess LLMs’ multi-step mental manipulation, we select a more complicated math magicbased on the Josephus Problem Schumer (2002). In this task, participants are asked to imagine four numbers and perform a sequence of operations, including duplication, rotation, and removal. The procedure is shown in Table 6. Ultimately, only two numbers remain, and mathematical constraints guarantee they are identical. In our experiment, LLMs are prompted to secretly think of four numbers and mentally execute the sequence of randomized operations. We measure the proportion of trials (out of 150) in which the LLM completes the task correctly.
Table 7: LLM performance on the math magic task.
@@
LLaMA-3.3-70B-Instruct-Turbo 7/150 5.7
LLaMA-3.1-8B-Instruct-Turbo 20/150 13.3
LLaMA-3.1-70B-Instruct-Turbo 7/150 5.7
LLaMA-3.1-405B-Instruct-Turbo 39/150 26.0
Qwen2.5-7B-Instruct-Turbo 8/150 5.3
Qwen2.5-72B-Instruct-Turbo 2/150 1.3
DeepSeek-V3 4/150 2.7
(a) LLMs without CoT. GPT-4.1-2025-04-14 fails to complete most of the cases, incorrectly assuming that the necessary numerical inputs are missing.
@@
GPT-4o-Mini-2024-07-18 5/150 3.3
GPT-4o-2024-05-13 26/150 17.3
GPT-4o-2024-08-06 31/150 20.7
GPT-4o-2024-11-20 - -
LLaMA-3.3-70B-Instruct-Turbo 25/150 16.7
Qwen2.5-7B-Instruct-Turbo 49/150 32.7
Qwen2.5-72B-Instruct-Turbo 37/150 24.7
DeepSeek-V3 48/150 32.0
o1-Mini-2024-09-12 75/150 50.0
o3-Mini-2025-01-31 145/150 96.7
o4-Mini-2025-04-16 54/150 36.0
QwQ-32B 135/150 90.0
DeepSeek-R1 150/150 100
(b) LLMs with CoT and Large Reasoning Models. GPT-4o-2024-11-20 consistently fails this task.
@@
Table 7a reports the accuracy of prompting models to output the two numbers directly. Most LLMs perform poorly on this task, with notably higher accuracy observed in the LLaMA model family. This finding aligns with results from the number guessing game shown in Table 1, where LLaMA models generate more realistic distributions than other models. Together, these consistent results suggest that current LLMs generally lack the working memory required for such tasks.
We further examine whether CoT prompting improves performance on this task. Table 7b presents the results of prompting models to reason step by step, as well as the performance of o1-like long reasoning models (LRMs). Base models prompted to reason step-by-step achieve 10–30% accuracy—substantially higher than without CoT. DeepSeek-R1 attains 100% accuracy, and other LRMs also perform well, indicating that extended token-level reasoning substantially improves accuracy. However, consistent with our number-guessing experiment, models exhibit a strong preference for the number seven, which appears to influence performance—especially within the o-series. Among o1-Mini’s 75 correct predictions, 50 (66.7%) select 7; for o3-Mini and o4-Mini, the counts are 68/145 (46.9%) and 37/54 (68.5%), respectively. These findings suggest that models may not fully solve the task but instead exploit biased number preferences: notably, o3-Mini—being least likely to guess 7—outperforms the other two o-series models. In conclusion, current LLMs struggle with a class of tasks demanding internal object imagination and mental manipulation, relying on an effective working memory.
@@ 6 Conclusion
The Use of Large Language Models
LLMs were employed in a limited capacity for writing optimization. Specifically, the authors provided their own draft text to the LLM, which in turn suggested improvements such as corrections of grammatical errors, clearer phrasing, and removal of non-academic expressions. LLMs were also used to inspire possible titles for the paper. While the system provided suggestions, the final title was decided and refined by the authors and is not directly taken from any single LLM output. In addition, LLMs were used as coding assistants during the implementation phase. They provided code completion and debugging suggestions, but all final implementations, experimental design, and validation were carried out and verified by the authors. Importantly, LLMs were NOT used for generating research ideas, designing experiments, or searching and reviewing related work. All conceptual contributions and experimental designs were fully conceived and executed by the authors.